Find the cube root of each of the following numbers by prime factorisation method.

 

(v) 15625

 

Grade:12th pass

1 Answers

Harshit Singh
2 years ago
Dear Student


(v) 15625

Solution:

15625 = 5×5×5×5×5×5

By grouping the factors in triplets of equal factors, 15625 = (5×5×5)×(5×5×5)

Here, 15625 can be grouped into triplets of equal factors,

∴ 15625 = (5×5) = 25

Hence, 25 is cube root of 15625.


Thanks